OpenStudy (anonymous):

Fill in the box with a whole number to make a true statement. little 3 little 9 8 3 = ? 9

OpenStudy (anonymous):



OpenStudy (anonymous):

@K12awesomeness

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@usercode3rror

OpenStudy (anonymous):

somebody explain this to me

jhonyy9 (jhonyy9):

do you think it 8^3 = x^9 find x = ?

OpenStudy (anonymous):

I'm not sure

jhonyy9 (jhonyy9):

but why you have wrote this question here in history column ?

OpenStudy (anonymous):

It changed on me sorry it was math

jhonyy9 (jhonyy9):

x^9 = (x^3)^3 8^3 = (x^3)^3 make cuberoot from both sides and will get 8= x^3 8=2^3 2^3 = x^3 make cuberoot from both sides and will get 2=x

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

np was my pleasure bye bye